Apple0291:

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: (2 points) 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

jhonyy9:

given that a mixture contain 1000 pounds of Arabica beans so us this value substitute it inside equation and so in this way try calcule the value of R

jhonyy9:

do you understand me ? please collaborate

Apple0291:

that all the test gave me and i am so lost on how to slove it

jhonyy9:

do you see above given this equation 700A + 1200R = 1,000,000 and you know that the value of A = 1000 so just substitute this 1000 in place of A inside equation and from this calcule the value of R

Apple0291:

what i am still lost

jhonyy9:

700A + 1200R = 1,000,000 please rewrite this equation so that in place of A write 1000

jhonyy9:

@jorellcross do you understand me ?

Apple0291:

100a+1200r=1000,000

jhonyy9:

yes this is the equation what you need using but for you can calcule the value of R you need substitute in place of A this given value what is 1000

jhonyy9:

no sorry you ve wrote wrong this equation here is correct 700A + 1200R = 1,000,000

jhonyy9:

what you get than in place of A write 1000 ?

Apple0291:

you got me even more lost

jhonyy9:

@jorellcross ?

jhonyy9:

700A + 1200R = 1,000,000 so substitute this given value of A in place of A and will get 700*1000 +1200R = 1000000 yes ?

jorellcross:

ohhh boy

jhonyy9:

and now how you get the value of R ?

jhonyy9:

@Apple0291 700*1000 = ?

jhonyy9:

please !!!

Apple0291:

-300

jorellcross:

here

jhonyy9:

* this sign mean time multiplie

jorellcross:

bommm bata bing

jhonyy9:

@jorellcross very very nice job CONGRATS !!!

jorellcross:

but she give me a medal

jorellcross:

not